OpenStudy (anonymous):

can someone show me the steps to the answer, i have the answer but need to see it written out. 4/x=9/5-7x-4/5x

OpenStudy (anonymous):

what answer do you have?

OpenStudy (anonymous):

answer is 8

OpenStudy (anonymous):

How did you get your answer without doing any steps?

OpenStudy (anonymous):

i just happened to guess the right answer because i didnt know it

OpenStudy (anonymous):

\[\frac{ 4 }{ x }=\frac{ 9 }{ 5 }-\frac{ 7x-4 }{ 5x }\]

OpenStudy (anonymous):

Is that the problem you were given? You are just restating it clearly?

OpenStudy (anonymous):

yes, so its easier to understand

OpenStudy (anonymous):

that's a good idea... the problem wasn't clear when typed out as text.

OpenStudy (anonymous):

@sue101 , you can explain if you'd like...

OpenStudy (anonymous):

i know, but it wont let me input the correct equation initially

OpenStudy (anonymous):

i know... no worries... just saying it was good to clarify.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you get a common denominator on the right side of the equation?

OpenStudy (anonymous):

oh go ahead

OpenStudy (anonymous):

Here's what I would do... On the right side of the equation, get a common denominator... \[\frac{ 9 }{ 5 } - \frac{ 7x-4 }{ 5x }\]

OpenStudy (anonymous):

is it 5 ?

OpenStudy (anonymous):

then multiply the first term by (x/x) to get 9x/5x Then you can find 9x/5x - (7x-4)/5x = [ 9x - (7x-4) ] / 5x simplify to (2x +4)/5x That is the right side of the equation.

OpenStudy (anonymous):

ahh, making more sense

OpenStudy (anonymous):

so overall... 4/x = (2x+4)/5x Multiply both sides by x: (4/x)(x) = (x)(2x+4)/5x Cancel the x's on the left and right: 4 = (2x+4)/5 Then multiply both sides by 5: 4(5) = (5)(2x+4)/5 = 2x+4 20 = 2x + 4 Subtract 4 from both sides: 20-4 = 2x 16 = 2x Divide both sides by 2 8 = x so x = 8

OpenStudy (anonymous):

when you multiply the x's, why do they cancelout

OpenStudy (anonymous):

nevermind, i actually get it :-)

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

Glad to help... so you understand the process pretty well now? Think you can take a shot at the next one? It's ok to read my solution, but you really learn it when you try working through it first on your own :) Good luck!